If you set the true probability of heads to ,0.5, what should be the true probability of tails? true probability of tails: How many heads would you expect to observe after 10 tosses? number of heads: Perform this experiment. Set the probability of heads to,0.5, and observe how many heads you get after 10 tosses. number of heads after 10 tosses: If the true probability of heads is 0.5, how many heads would you expect to observe after 50 tosses? number of heads after 50 tosses: If the true probability of heads is 0.5, how many heads would you expect to observe after 500 tosses? number of heads after 500 tosses:
Answer #1

True probability of tails = 1 - True probability of heads

= 1 - 0.5 = 0.5

Number of heads expected after 10 tosses = 0.5*10 = 5

Number of heads observed after 10 tosses = 6

Number of heads expected to observe after 50 tosses = 50*0.5 = 25

Number of heads expected to observe after 500 tosses = 500*0.5 = 250